OpenStudy (anonymous):

if log2 = 0.3010 and log3 = 0.4771; then log300 - log8 = 1.5741 .... true or false? why?

OpenStudy (anonymous):

but log5 isnt given...ur only supposd to use log2 and log3

OpenStudy (bahrom7893):

log(300)=log(3*100)=log(3*10^2)=log3 + 2log10

OpenStudy (bahrom7893):

Sorry I was supposed to stop there

OpenStudy (bahrom7893):

=log3+2*1=2+log3

OpenStudy (bahrom7893):

log8=log2^3=3log2

OpenStudy (bahrom7893):

So just find: 2+log3-3log2

OpenStudy (bahrom7893):

And that does in fact equal 1.5741

OpenStudy (anonymous):

log300 - log8 \[\log(3\times100)-\log2^{3}=\log3+2\log10-3\log2=\log3+2\times1-3\log2\] =0.4771 +2-3*0.3010 =2.4771 -0.9030=1.5741 Thus the answer is true

OpenStudy (anonymous):

thank u :)